OpenStudy (jenniferjuice):

Choose the correct simplification of the expression (3x − 6)(2x^2 − 4x − 5) 6x^3 − 24x^2 + 9x − 30 6x^3 + 9x + 30 6x^3 − 24x^2 + 9x + 30 6x^3 − 24x^2 + 39x + 30

OpenStudy (uri):

Use the foil method (3x − 6)(2x^2 − 4x − 5) 3x(2x^2-4x-5)-6(2x^2-4x-5) Solve this and tell me what you get.

OpenStudy (uri):

@jenniferjuice

OpenStudy (jenniferjuice):

wait if i were to use the foil methos 3x(2x^2 - 4x - 5) - 6 (2x^2 - 4x - 5) first would i multiply 3x and -6 ?

OpenStudy (uri):

You have to mutiply 3x with with whole second bracket and -6 with the whole 2nd bracket 3x(2x^2 - 4x - 5) What do you get? @jenniferjuice

OpenStudy (jenniferjuice):

so .... -18x

OpenStudy (jenniferjuice):

and 3x * -5

OpenStudy (uri):

3x*2x^2=6x^3 3x*-4x=-12x^2 3x*-5=15x

OpenStudy (uri):

Now multiply this 6 (2x^2 - 4x - 5) -6*+2x^2=-12x^2 -6*-4x=-28x 6*-5=-30

OpenStudy (jenniferjuice):

i got -18x - 15x + 30 +25

OpenStudy (uri):

So, 3x(2x^2 - 4x - 5) - 6 (2x^2 - 4x - 5) 6x^3-12x^2-15x-12x^2+24x+30

OpenStudy (uri):

Now bring all x^3 values together,all x^2 values together and all x values together and numericals together